you are provided with the following information economy

Y=C+I+G+X-M

C=200+0.8Yd

I0=500

G0=500

X0=100

M=50+0.4Y

using the data provided demonstrate that multiplier for open economy is smaller than the multiplier for closed economy


1
Expert's answer
2021-09-19T18:32:19-0400

Let's T=0.100

for open economy:

Y=C+I+G+X-M

C=200+0.8Yd=200+0.8(Y-T)=200+0.8(Y-100)

I0=500

G0=500

X0=100

M=50+0.4Y

Y=200+0.8(Y-100)+500+500+100-(50+0.4Y)=200+0.8Y-80+500+500+100-50-0.4Y=0.4Y+1170

Y-0.4Y=1170

0.6Y=1170

Y=1950


for closed economy

Y=C+I+G=200+0.8(Y-100)+500+500=200+0.8Y-80+500+500=0.8Y+1120

Y-0.8Y=1120

0.2Y=1120

Y=5600

According to the Keynesian theory, an increase in consumption, government spending or investment leads to an increase in national income (total output) , and this growth will be greater than the growth of any part of spending.

The value of the multiplier calculated for an open economy will be less than for a closed one. This is due to the fact that the costs of importing products, as well as savings, cease to be a component of the total demand for the domestic product.

The expenditure multiplier shows the geometric progression of the increase in initial expenditures to the economy, that is, an increase in expenditures (investments) causes a greater increase in national income than the initial amount of expenditures.